Circular No. 6923

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 K5 (LINEAR) It appears that yet another comet has been recognized among the fast-

moving objects reported by the Lincoln Laboratory Near-Earth Asteroid Research project and included on The NEO Confirmation Page:

1998 UT  α_{2000}  δ_{2000}  m_{2}
May 26.37939  17^{h}26^{m}02.^{s}33  +2^{o}02′11′′.1  18.0

The high orbital eccentricity indicated in the early computations was the first clue that the object might be a comet, although there were the complications that the orbit has a low inclination and that, when asked, several observers felt that their data showed the object to be asteroidal in nature. The first observational indication of cometary character was provided by P. J. Shelus, who remarked that images obtained by J. G. Ries (McDonald, 0.76-m reflector, poor seeing) on May 29.39 UT seemed to show a tail to the south and west, roughly opposite the object’s direction of motion. P. Pravec (Ondřejov, 0.65-m reflector) also remarked on the tail, consistently on coadded images and marginally on individual images, 0′.4 long in p.a. 235^{o} on May 27.03, 0′.4 long in p.a. 225^{o} on May 28.95 and 0′.3 long in p.a. 227^{o} on May 29.97. On May 30.46 C. Veillet (Canada-France-Hawaii, 3.6-m reflector) noted a clear 30′′ tail on a 1-min exposure

(http://www.cfht.hawaii.edu/~veillet/A65.html). S. Nakano reports

that A. Sugie (Dynic, 0.6-m reflector, bad seeing), observing on May 27.58, independently noted (on four coadded frames) a 16′′ tail in p.a. 230^{o}.

The following parabolic orbital elements, computed by the undersigned from 128 observations May 26–30, are taken from MPEC 1998-K33:

T = 1998 July 17.326 TT  ω = 99.^{o}215
Ω = 211.333 }2000.0
q = 0.96411 AU  i = 10.044
1998TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
May 31 17^{h}41.^{m}26 + 7^{o}32.^{′}6 0.283 1.258 145.^{o}9 26.^{o}9 15.3
June 2 17 49.25 +10 21.5 0.266 1.237 143.2 29.4 15.0
4 17 58.33 +13 27.8 0.250 1.217 140.1 32.3 14.8
6 18 08.72 +16 52.4 0.236 1.197 136.6 35.6 14.6
8 18 20.68 +20 35.1 0.223 1.178 132.7 39.3 14.5
10 18 34.53 +24 34.4 0.212 1.159 128.4 43.4 14.3
12 18 50.64 +28 47.2 0.203 1.141 123.7 47.8 14.1
14 19 09.40 +33 07.6 0.196 1.123 118.7 52.5 14.0
16 19 31.22 +37 27.4 0.191 1.106 113.5 57.4 13.8
1998 May 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden